OpenStudy (anonymous):

(2+sqrt-16)(4-sqrt-9) Do the expression and express the answer in a+bi form.

OpenStudy (alex_mattucci):

could you draw this one out please?

OpenStudy (anonymous):

\[(2+\sqrt{-16})(4-\sqrt{-9})\]

OpenStudy (alex_mattucci):

ok thank you

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \left(2+\sqrt{-16}\right) \left(4-\sqrt{-9}\right) }\) you know that: \(\sqrt{-a^2}=\sqrt{a^2}\times\sqrt{-1}=ia\) And this way, it follows that: \(\bullet~~~~~~\sqrt{-9}=3i \) \(\bullet~~~~~~\sqrt{-16}=4i\)

OpenStudy (solomonzelman):

So, you can re-write your expression is: \(\large\color{black}{ \displaystyle \left(2+4i\right) \left(4-3i\right) }\) then, expand the expression: \(\large\color{black}{ \displaystyle 2\left(4-3i\right)+4i\left(4-3i\right) }\) continue expanding, and tell me what you get:

OpenStudy (anonymous):

Before I continue, why did you put a 2 and 4i in front of the binomials?

OpenStudy (solomonzelman):

Have you expanded expressions like: (A+B)(C+D) previously?

OpenStudy (solomonzelman):

((You would multiply times A times C+D, and B times C+D. Then you add the results.))

OpenStudy (anonymous):

yes but I don't understand how you did that

OpenStudy (anonymous):

Oh!

OpenStudy (solomonzelman):

Yes?

OpenStudy (solomonzelman):

So can you continue from: \(\large\color{black}{ \displaystyle 2\left(4-3i\right)+4i\left(4-3i\right) }\) for me please?

OpenStudy (anonymous):

Yes Im working it out now

OpenStudy (anonymous):

I got a trinomial?

OpenStudy (solomonzelman):

i is not a variable, it is a square root of -1. And it abides by the property: \(i^2=\left(\sqrt{-1}\right)^2=-1\)

OpenStudy (solomonzelman):

So the term of i² is just a negative one.

OpenStudy (solomonzelman):

Anyway, can you post what you have got?

OpenStudy (anonymous):

when I work it out I get 8+22i which isn't the answer

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2\left(4-3i\right)+4i\left(4-3i\right) }\) \(\large\color{black}{ \displaystyle 2(4)+2(-3i)+4i(4)+4i(-3i) }\) \(\large\color{black}{ \displaystyle 8+(-6i)+(8i)+(-12i^2) }\)

OpenStudy (anonymous):

4i(4) should be 16i why is it 8i?

OpenStudy (anonymous):

This is where I got my answer wrong

OpenStudy (solomonzelman):

Oh, sorry, my bad

OpenStudy (solomonzelman):

yes, you were right.... \(\large\color{black}{ \displaystyle 8+(-6i)+(16i)+(-12i^2) }\) \(\large\color{black}{ \displaystyle 8+2i+(-12(-1)) }\) \(\large\color{black}{ \displaystyle 8+10i+(12) }\) \(\large\color{black}{ \displaystyle 20+10i }\)

OpenStudy (anonymous):

Wow I made a really dumb mistake.. Thanks!

OpenStudy (solomonzelman):

I did as well.